OpenStudy (anonymous):

Please help! (: Been stuck on this problem for 2 hrs now. A particle is moving along the curve y=3(sqrt(4x+4)). As the particle passes through the point (3, 12), its x-coordinate increases at a rate of 4 units per second. Find the rate of change of the distance from the particle to the origin at this instant.

OpenStudy (anonymous):

looks like you need to set up parametric equations, because this deals with 3 variables: x, y, and t (time)

OpenStudy (anonymous):

How do I go about doing that. I started with \[d ^{2}=x^{2}+9(4x+4)\] and got \[d^{2}=x^{2}+36x+36\] then I used the points (0,0) and (3,12) to find the distance. but I don't know where to go from there.

OpenStudy (anonymous):

I got 2dd'=2xx'+6x' =2(4)(4)+6(4) =56

OpenStudy (anonymous):

^what's that?that ain't integration

OpenStudy (anonymous):

oh my bad..i had to say derivative..GOSH why did i say that

OpenStudy (anonymous):

first of all differentiate this one \[d ^{2}=x ^{2}+36x+36\] @pdd21

OpenStudy (whpalmer4):

@pdd21 That answer doesn't look plausible. At t = 0, the particle is at (3,12). 1 second later, it's at \((7,3\sqrt{4*7+4})\) or \((7,12\sqrt{2})\). If you take the distance between those points, \[d = \sqrt{(7-3)^2+(12\sqrt{2}-12)^2} \approx 6.4\]

OpenStudy (anonymous):

wouldn't it be 2x+36? @stgreen

OpenStudy (anonymous):

2dd'=2x+36 where d'=rate of change of distance w.r.t x-coordinate=4

OpenStudy (anonymous):

and d would be \[\sqrt{153}\] right? @stgreen

OpenStudy (anonymous):

what?why?

OpenStudy (whpalmer4):

yes, \(3\sqrt{17}\) is distance from origin to (3,12)

OpenStudy (anonymous):

oh yeah right

OpenStudy (anonymous):

I used the distance formula to find d? sorry, I suck at math. @stgreen

OpenStudy (anonymous):

So now that I have 2dd'=2x+36 and the distance where do I go from there?

OpenStudy (anonymous):

if you have distance d..you should know x too

OpenStudy (whpalmer4):

\[d = \sqrt{(x-0)^2+(3\sqrt{4x+4} - 0)^2} = \sqrt{x^2+9(4x+4)} = \sqrt{x^2+36x+36}\]\[\frac{d}{dx}[(x^2+36x+36)^{1/2} = \frac{1}{2}*(2x+36)(x^2+36x+36)^{-1/2}\]Evaluate that at \(x=3\), multiply by \(dx/dt = 4\) and you've got your answer....

OpenStudy (anonymous):

^yeah right

OpenStudy (anonymous):

The way you layed out the problem made so much more sense. I see where I went wrong.. I derived wrong. Thank you so much! I got 6.79 and it was correct! @whpalmer4

OpenStudy (whpalmer4):

You're welcome. I knew there was something wrong with what you were doing, but I hadn't figured out just what!